# PersonaCraft: Personalized Full-Body Image Synthesis for Multiple Identities from Single References Using 3D-Model-Conditioned Diffusion

### v1

@@ 5 Discussion
While our method is versatile and can be applied to other ControlNet models, the performance of our face personalization depends significantly on the underlying face network. Additionally, the accuracy of 3D human model fitting is contingent upon the performance of the fitting algorithm used. Variations in the quality of the fitting process may impact the final output, especially in cases where the reference data is incomplete or inaccurate. It is crucial to emphasize that PersonaCraft must be used responsibly, as the misuse of such technology for malicious purposes, including unauthorized identity manipulation, poses ethical and legal concerns.
@@ 6 Conclusion
In this paper, we presented PersonaCraft, a powerful framework for generating highly personalized images of multiple individuals, even in complex occlusion scenarios. By integrating a 3D human modeling approach with SMPLx-ControlNet, PersonaCraft effectively handles occlusions and ensures precise full-body personalization—advancing beyond previous methods that primarily focused on face identity or struggled with overlapping poses. The framework also introduces a novel feature for user-defined body shape control, which allows for more tailored and flexible customization, making it applicable to a wide range of scenarios.
Through comprehensive experimentation, PersonaCraft demonstrated superior performance in preserving individual identities, handling intricate poses, and maintaining body shape consistency across multiple individuals in a scene. We believe PersonaCraft holds significant promise for enhancing personalized content creation in diverse domains, including virtual environments, social media, and more. Future work will focus on refining the system to handle more complex interactions and real-world applications for dynamic scene personalization.
Supplementary Material
@@ Appendix A Additional Results
A.1 Additional Qualitative Comparison
As shown in Fig. S1, our method significantly outperforms existing approaches. InstantID+OMG[71, 43] and IP-Adapter+OMG [80, 43], which use 2D skeleton-based pose conditioning, show severe anatomical inaccuracies in challenging cases (yellow arrows) due to the limitations of 2D pose representations. Blue arrows highlight that our method preserves body structure and pose fidelity while excelling in face identity and body shape consistency.
DreamBooth [60], lacking pose guidance, exhibits worse distortions, omitting individuals in multi-person scenes and misaligning clothing with body shapes.
These results highlight the robustness of PersonaCraft as a state-of-the-art solution for personalized image generation in complex scenarios.
A.2 Comparison with Additional Baselines
We compared PersonaCraft against additional baselines, including UniPortrait [32], MS-Diffusion [73], and FastComposer [76]. While these methods share similar capabilities, they are not fully suited for our benchmark, making direct comparisons challenging.
As shown in Fig. S2 and S3, yellow arrows highlight anatomical inconsistencies in complex poses and occluded scenarios due to reliance on 2D pose representations or the absence of pose control. PersonaCraft, in contrast, generates anatomically accurate and natural images under these conditions.
Blue arrows mark individuals used to evaluate body shape preservation, where PersonaCraft outperforms baselines, which frequently fail to maintain accurate body shapes.
Pink arrows indicate the inability of UniPortrait and FastComposer to handle multiple identities effectively, often mixing or duplicating identities. PersonaCraft successfully preserves distinct face identities for all individuals.
@@ A.2 Comparison with Additional Baselines
A.3 Versatility of SCNet
@@
A.4 Ablation Study on Conditioning Scale
@@
A.5 Additional Results of User Study
We present the cumulative percentage of the Top 3 results in Tab. S1, with details on the user study settings provided in the supplementary Sec. B.4. Consistent with the Top 1 results in Tab. 2 of the main paper, our method is preferred over other baselines in terms of all attributes.
Table S1: The cumulative percentage of Top 3 from user study on naturalness, face identity, body shape preservation, and text-image correspondence.
[…]
(Continued in Sec. B)
@@ A.5 Additional Results of User Study
Figure S2: Comparison of PersonaCraft with UniPortrait [32], MS-Diffusion [73], and FastComposer [76]. Yellow arrows show anatomical inconsistencies in poses and occlusions. Pink arrows highlight identity mixing or duplication in baselines. MS-Diffusion copies clothing without proper adjustment.PersonaCraft excels in face identity, body shape consistency, and naturalness compared to the baselines.
Figure S3: Comparison of PersonaCraft with UniPortrait [32], MS-Diffusion [73], and FastComposer [76]. Yellow arrows show anatomical inconsistencies in poses and occlusions. Pink arrows highlight identity mixing or duplication in baselines. Blue arrows refer to the people for checking correct body shape preservation. MS-Diffusion copies clothing without proper adjustment.PersonaCraft excels in face identity, body shape consistency, and naturalness compared to the baselines.
@@ B.1 Additional Implementation Details
Face and body identity extraction. We selected MultiHMR [12] as our SMPLx fitting method. MultiHMR is a single-shot model that recovers 3D human meshes from a single RGB image, leveraging the SMPLx parametric model to predict full-body meshes, including hands and facial expressions, with 3D localization in the camera coordinate system. The body shape parameters, $\bm{\beta}$, are 10-dimensional vectors multiplied element-wise with orthonormal shape displacement components.
For facial identity processing, we use the antelopev2 facial detection and recognition models from InsightFace [1] to detect and extract 512-dimensional face identity embeddings, $\bm{f}$, from human images.
3D-aware pose conditioning using SMPLx-ControlNet. We base our SCNet on controlnet-union-sdxl-1.0 [77] and fine-tune it for SMPLx [57] depth conditioning. The architecture supports more than 10 control types for high-resolution text-to-image generation, with depth selected as the control type in our implementation.
@@
Text captions for MPII-SMPLx were derived using GPT-4 [3] by Khan et al. [38], providing detailed descriptions of samples from the MPII dataset.
Multiple human personalized image synthesis. We used the YamerMIX-v8 version of SDXL to enhance the visual quality of human-centric scenes. We selected IdentityNet from InstantID [71] as our face identity personalization module, which enables instant, zero-shot, identity-preserving image generation. The IdentityNet design imposes strong semantic and weak spatial conditions, integrating facial and landmark images with textual prompts to guide the image generation. For face landmarks, following InstantID, we used five key facial points (two for the eyes, one for the nose, and two for the mouth) as spatial control signals, applying a more generalized constraint compared to detailed key points.
@@ B.2 Details on Metrics
Figure S6: Example images from the user study.
@@ B.3 Details on Baselines
To evaluate PersonaCraft, we compared it with several baselines for single-shot, multi-identity, and pose-controllable image synthesis, all implemented using SDXL [59]. Key baselines include OMG [43] for multi-concept personalization, InstantID/IP-Adapter [71, 80] for single-shot personalization, 2D pose ControlNet [81], and optimization-based methods like DreamBooth [60] and Texture Inversion [23].
OMG + InstantID/IP-Adapter/IPA-Face. OMG [43] introduces a two-stage sampling solution for multi-concept personalization. The first stage handles layout generation and occlusion management, while the second stage integrates concepts using visual comprehension and noise blending. OMG can also be combined with single-concept models like InstantID without additional tuning. For OMG+InstantID, we follow the official inference code from the InstantID repository [44]. For OMG+InstantID and OMG+IP-Adapter/IPA-Face, we replace InstantID with IP-Adapter and IPA-Face, respectively, to adapt the framework for different face identity modules.
@@ B.4 Details on User study
We conducted a user study to assess the naturalness, face identity preservation, body shape preservation, and text-image correspondence of images generated by three baseline methods (one from each group) and our method. Participants ranked the top three methods based on the following criteria: 1) Naturalness: Rank the images in order of the most natural-looking, considering factors such as physically impossible appearances, illogical features, inconsistencies, or lack of real-world physics and connections. 2) Face Identity Preservation: Rank the images in order of how well they reflect the face identity of the personalized character. 3) Body Shape Personalization: Rank the images based on how accurately they reflect the personalized character’s body shape. 4) Text Correspondence: Rank the images based on how closely they align with the textual description. The study collected a total of 10,332 responses from 123 participants across 7 cases, including both custom and COCO-Wholebody scenarios. We present illustrative example images from the user study in Fig. S6.
Acknowledgement
This work is supported by Samsung Electronics MX Division.

### v2

@@ 5 Limitations
While our method is versatile and can be applied to other ControlNet models, the performance of our face personalization depends significantly on the underlying face identity network. Additionally, the accuracy of 3D human model fitting is dependent on the performance of the fitting algorithm used. Variations in the quality of the fitting process may impact the final output, especially in cases where the reference data is incomplete or inaccurate. Failure cases can be found in the supplementary material.
@@ 6 Conclusion
We present PersonaCraft, a framework for occlusion-robust, full-body personalized image synthesis in complex multi-human scenes. PersonaCraft integrates diffusion models with 3D human modeling, using SMPLx-ControlNet for enhanced pose conditioning. To address occlusions at a fine level, we introduce Occlusion Boundary Enhancer and Occlusion-Aware Classifier-Free Guidance strategy, improving conditioning in occluded areas. Seamlessly combined with Face Identity ControlNet, our approach enables full-body personalization beyond facial identity. Extensive experiments show PersonaCraft outperforms prior arts in personalization accuracy and occlusion robustness in complex scenes.
@@ Appendix A Additional Results
A.1 Personalized Multi-Human Scene Generation
Additional Qualitative Comparison. As shown in Fig. S13 and  S14, our proposed method demonstrates significant advantages over existing approaches. Notably, methods like InstantID+OMG [81, 43] and IPAdapter+OMG [90, 43], which rely on 2D skeleton-based pose conditioning, exhibit severe anatomical inaccuracies in challenging scenarios (highlighted by yellow arrows). These issues stem from the inherent limitations of 2D pose representations, which struggle to handle overlapping body parts and intricate interactions effectively. Blue arrows highlight cases for verifying correct body shape preservation, where our approach maintains accurate body structure and pose fidelity while delivering superior performance in both face identity preservation and body shape consistency.
DreamBooth [66], on the other hand, suffers even more pronounced issues due to its lack of pose guidance. This leads to severe anatomical distortions and, in some cases, the complete omission of individuals in multi-person scenes. Additionally, DreamBooth struggles with clothing-body shape displacement, where clothing styles are directly transferred without adapting to the individual’s body shape.
These findings further underscore the robustness and versatility of PersonaCraft, making it a state-of-the-art solution for personalized image generation in complex, real-world scenarios.
Comparison with Additional Baselines. We compared PersonaCraft against other baselines, including, UniPortrait [31], MS-Diffusion [83], and FastComposer [86]. While these methods share similar capabilities, they are not fully suited for our benchmark, making direct comparisons challenging. As shown in (Fig.S15), yellow arrows highlight anatomical inconsistencies in complex poses and occluded scenarios due to reliance on 2D pose representations or the absence of pose control. PersonaCraft, in contrast, generates anatomically accurate and natural images under these conditions.
@@ A.1 Personalized Multi-Human Scene Generation
A.2 Pose-Controlled Multi-Human Scene Generation
We assess pose-controlled multi-human scene generation by comparing PersonaCraft with existing methods in both single-human and multi-human contexts. As shown in Fig. S16S17, and S18, our proposed method demonstrates significant advantages over existing approaches. Notably, methods that rely on 2D pose conditioning (e.g., T2IAdapter-SDXL[56], ControlNet-SDXL [91, 64]) struggle with occlusion handling, resulting in misaligned poses and distorted structures. ControlNet-Flux [48, 91] shows instability: at a lower conditioning scale (0.8), it fails to preserve anatomical coherence, while at a higher scale (1.0), pose accuracy improves but image fidelity decreases. In contrast, PersonaCraft successfully balances pose accuracy, body realism, and image fidelity, highlighting the effectiveness of our 3D-aware pose conditioning.
A.3 PersonaCraft with Stylization
The proposed method is a plug-and-play approach, making it compatible with various style-specific LoRAs. To evaluate its effectiveness, we conducted experiments combining PersonaCraft with diverse style LoRAs, including Crayon [59], Pastel [6], 3D Render [26], Pixel Art [57], Illustration [5], Frosting Lane [3], Pokémon Trainer [76], JoJo [58], Graphic Novel [15], and Cartoon [4]. The results, shown in Fig. S19, highlight the method’s ability to adapt to different styles effectively. Notably, styles such as Pastel, Illustration, JoJo, and Pokémon Trainer introduce changes in facial and body characteristics, occasionally altering perceived identity, due to their bias. Nevertheless, the outcomes remain visually compelling and demonstrate the versatility of our approach.
A.4 Versatility of SCNet
@@
A.5 Ablation Study on Conditioning Scale
@@
A.6 Ablation Study on Body Shape Parameters
For full-body personalized image generation, we extract the body shape parameters of the character to be personalized and use them for SMPLx rendering, which serves as the conditions for SCNet. In Tab. S6, we analyze the impact of incorporating the body shape parameters in this process. Using body shape parameters enhances body shape preservation during personalization. This indicates that leveraging the body shape parameters enables the generation of personalized images that more accurately reflect the character’s true physique.
Table S6: Evaluation of body shape preservation with and without the use of the body shape parameter.
Single Multi Total w/o body shape 0.615 0.520 0.539 w/ body shape 0.630 0.548 0.615
A.7 Ablation Study on Occlusion-aware 3D pose & Shape conditioning
Comparison of 3D Pose Representations for SCNet. In Fig. S22, we compare different combinations of SMPLx rendering-depth, normal, and RGB rendering-as conditioning inputs for SCNet. Using both depth and normal enables the model to leverage occlusion cues from depth and surface orientation information from normal. This leads to improved generation performance in occluded or complex body regions compared to using depth alone. However, incorporating RGB rendering in addition to depth and normal degrades image quality due to the use of multiple ControlNets for the same region. Therefore, we adopt the combination of depth and normal as base conditioning for SCNet.
Effectiveness of OccNet and OccCFG. Also, the effect of our occlusion-aware 3D pose & shape conditioning components is analyzed in Fig. S22. Using only 2D pose leads to structural inconsistencies, while relying solely on 3D pose (SCNet) struggles with fine-grained occluded regions. Our full model (SCNet + OccNet + OccCFG) effectively preserves pose structure while handling occlusions.
A.8 Efficiency Analysis
As analyzed in Tab. S7, we compare the inference times for multi-identity personalized synthesis across different methods, specifically for generating images with three distinct identities. Textual Inversion [22] and DreamBooth [66], which rely on optimization-based personalization, require a batch size of 4 and 500 optimization steps per identity. This process results in significantly longer inference times, making these methods highly inefficient for real-time applications.
Table S7: Inference times for multi-identity personalized synthesis.
Method Total Time (secs) Text Inversion 1636.15 DreamBooth 770.713 InstantID + OMG 46.94 IPAdapter + OMG 44.62 IPA-Face + OMG 35.46 PersonaCraft (ours) 17.25
A.9 Additional Results with Baselines Finetuned on our Training Dataset
We further compare PersonaCraft with key baselines finetuned on our training dataset: InstantID + OMG, IPAdapter + OMG, and IPA-Face + OMG for personalized multi-human scene generation, and ControlNet-SDXL for pose-controlled multi-human scene generation.
Personalized Multi-Human Scene Generation.
Table S8: Additional comparison of baseline models fine-tuned on our training dataset (MPII). Quantitative evaluation of multi-human personalization across face identity, body shape preservation, pose accuracy, text alignment, and image quality. ‘Single’ denotes personalization for a single individual, while ‘Multi’ refers to cases with multiple identities (2–5), with ‘Total’ representing the averaged results. (*: fine-tuned on our training dataset.)
Multi-Human Personalization Face ID preservation↑ Body shape preservation↑ Pose Text Image quality Single Multi Total Single Multi Total MPJPE (3D)↓ AP-0.5 (2D)↑ CLIP sim ↑ IS ↑ KID ↓ InstantID + OMG 0.418 0.220 0.252 0.563 0.427 0.448 85.624 0.333 0.264 3.809 0.0996 IPAdapter+ OMG 0.204 0.155 0.162 0.606 0.45 0.472 84.893 0.362 0.267 3.736 0.0984 IPA-Face + OMG 0.350 0.181 0.207 0.568 0.429 0.451 86.373 0.355 0.265 3.930 0.0974 Ours 0.421 0.298 0.317 0.630 0.548 0.560 60.654 0.506 0.273 4.238 0.0931
Although the baselines are fine-tuned on our training dataset, Tab. S8 shows that our method consistently outperforms baselines in identity and body shape preservation. It achieves the lowest MPJPE (3D) and highest AP (2D), indicating superior alignment with input poses. Additionally, PersonaCraft surpasses baselines in IS and KID, demonstrating enhanced perceptual quality and text-image coherence.
Pose-Controlled Multi-Human Scene Generation.
Table S9: Additional comparison of the baseline fine-tuned on our training dataset (MPII). Quantitative evaluation of pose-controlled human generation in terms of pose accuracy, text alignment, and image quality. (*: fine-tuned on our training dataset.)
Pose-Controlled Human Generation Pose Text Image quality MPJPE (3D) ↓ AP-0.5 (2D) ↑ CLIP sim ↑ IS ↑ KID ↓ ControlNet-SDXL 100.817 0.313 0.281 3.407 0.122 Ours 62.647 0.495 0.274 4.114 0.091
Although the baselines are fine-tuned on our training dataset, Tab. S9 shows that PersonaCraft achieves the lowest MPJPE (3D) and highest AP-0.5 (2D), demonstrating superior pose alignment and keypoint localization.
A.10 Failure Cases
While our method is versatile and can be applied to other ControlNet models, the performance of our face personalization depends significantly on the underlying face network. Additionally, the accuracy of 3D human model fitting is dependent on the performance of the fitting algorithm used. Variations in the quality of the fitting process may impact the final output, especially in cases where the reference data is incomplete or inaccurate as presented in Fig. S23.
(Additional Experimental details are continued in Sec. B)
@@ A.10 Failure Cases
Figure S14: Additional comparison of generated images. Yellow arrows highlight anatomical issues in InstantID+OMG [81, 43] and IPAdapter+OMG [90, 43] due to 2D pose limitations. DreamBooth [66] shows severe distortions and clothing mismatches. PersonaCraft excels in face identity, body shape consistency, and naturalness while being over twice as fast as OMG-based methods [81, 90, 43].
Figure S15: Comparison of PersonaCraft with UniPortrait [31], MS-Diffusion [83], and FastComposer [86]. Yellow arrows show anatomical inconsistencies in poses and occlusions. MS-Diffusion copies clothing without proper adjustment.PersonaCraft excels in face identity, body shape consistency, and naturalness compared to the baselines.
Figure S16: Qualitative comparison of 3D-aware pose conditioning for multi-human generation, covering both single and multi-human scenarios. PersonaCraft achieves superior alignment with the input pose while effectively handling occlusions, allowing for natural human anatomy to be maintained even in complex multi-human interactions. Our method outperforms baselines in preserving identity, body shape, and overall human realism. yellow arrows highlight unnatural anatomical structures.
Figure S17: Qualitative comparison of 3D-aware pose conditioning for multi-human generation, covering both single and multi-human scenarios. PersonaCraft achieves superior alignment with the input pose while effectively handling occlusions, allowing for natural human anatomy to be maintained even in complex multi-human interactions. Our method outperforms baselines in preserving identity, body shape, and overall human realism. yellow arrows highlight unnatural anatomical structures.
Figure S18: Qualitative comparison of 3D-aware pose conditioning for multi-human generation, covering both single and multi-human scenarios. PersonaCraft achieves superior alignment with the input pose while effectively handling occlusions, allowing for natural human anatomy to be maintained even in complex multi-human interactions. Our method outperforms baselines in preserving identity, body shape, and overall human realism. yellow arrows highlight unnatural anatomical structures.
Figure S19: Results of combining PersonaCraft with various style LoRAs, showcasing adaptability to styles like Pastel, JoJo, and Pokémon Trainer. Some styles alter facial and body identities due to their bias, while producing visually impressive outcomes.
@@
Figure S22: Ablation study on occlusion-aware 3D pose & shape conditioning. The combination of depth and normal as conditioning inputs for SCNet achieves the best generation performance in occluded or complex regions. While using SCNet faces issues preserving pose structure in fine-grained occluded regions, adding OccNet and OccCFG effectively addresses these problems.
Figure S23: Failure cases. The accuracy of 3D human model fitting depends on the fitting algorithm used, and variations in fitting quality, particularly when the reference data is incomplete or inaccurate, can impact the final output.
@@ B.1 Additional Implementation Details
Additional Details on Training Dataset. To account for occlusion scenarios, we balance the MPII [7] dataset with a 2:1 ratio of single-person to multi-person images. Depth clipping is applied during depth rendering to retain only values below 5, ensuring consistent quality. After preprocessing, we curate a final dataset of 6,348 image-text-SMPLx-parameter pairs. This carefully curated dataset enables robust model training with diverse 3D human poses, complex interactions, and detailed human parameters such as body shape and pose conditioning.
Details on Training of SCNet and OccNet. We base our SCNet on controlnet-union-sdxl-1.0 [87] and fine-tune it for SMPLx [61] depth-normal-ocluded edge conditioning. The architecture supports more than 10 control types for high-resolution text-to-image generation, with depth selected as the control type in our implementation.
@@
For depth edge extraction, we employed the Canny edge detector for more robust edge extraction instead of thresholding the spatial partial gradient of depth with $\tau$, using a low threshold of 5 and a high threshold of 15.
Details on Full-Body Personalized Image Synthesis. We adopt MultiHMR [11] as our SMPLx fitting method. MultiHMR is a single-shot model that reconstructs 3D human meshes from a single RGB image, leveraging the SMPLx parametric model to predict full-body meshes, including hands and facial expressions, with 3D localization in the camera coordinate system. The body shape parameters, $\bm{\beta}$, are represented as 10-dimensional vectors, each scaled by orthonormal shape displacement components.
For facial identity processing, we employ the antelopev2 facial detection and recognition models from InsightFace [1] to extract 512-dimensional face identity embeddings, $\bm{f}$, from human images.
To enhance the visual quality of human-centric scenes, we utilize the YamerMIX-v8 variant of SDXL. For face identity personalization, we incorporate IdentityNet from InstantID [81], which enables instant, zero-shot, identity-preserving image generation. IdentityNet enforces strong semantic and weak spatial conditions by integrating facial and landmark images with textual prompts to guide the generation process.
Following InstantID, we use five key facial landmarks (two for the eyes, one for the nose, and two for the mouth) as spatial control signals, providing a more generalized constraint than detailed key points.
Details on Dual-Pathway Body Shape Personalization. In this method, a CLIP [65] -based classifier is employed to extract body shape attributes in the form of text descriptions. These descriptions categorize the body type into various categories, such as "overweight," "muscular," "fat," etc. This is achieved by using a combination of CLIP, which bridges the gap between vision and text, and specific regional prompting techniques.
The body shape information is then used in Regional Diffusion, a concept derived from MultiDiffusion [10], where each diffusion timestep involves conditioning on both the body pose and a corresponding text description about the individual’s body shape. The process operates on each person’s instance throughout multiple diffusion timesteps, ensuring that the shape-specific features are incorporated and aligned with the pose dynamics.
Incorporating these shape attributes into the diffusion process allows the model to better represent personalized body shapes in the generation process, resulting in a more accurate and detailed synthesis of human shapes. This integration is achieved through the use of regional prompting, where at each timestep, the model is conditioned on both the body pose and the specific body shape description to refine the body shape in the generated instance. This process is further integrated with SCNet, a network that helps guide the shape refinement and personalization.
User-Defined Body Shape Control. PersonaCraft enables user-defined body shape control, allowing adjustments based on user preferences: 1) Reference-based control: The target body shape parameter, $\bm{\beta}_{\text{target}}$, is obtained from a reference image via SMPLx fitting. 2) Interpolation/extrapolation-based control: Given two reference body shape parameters, $\bm{\beta}_{1}$ and $\bm{\beta}_{2}$, the target shape is computed as $\bm{\beta}_{\text{target}}=\gamma\bm{\beta}_{1}+(1-\gamma)\bm{\beta}_{2}$, where $\gamma$ controls the interpolation/extrapolation ratio. The resulting $\bm{\beta}_{\text{target}}$
@@ B.2 Details on Metrics
Figure S24: Example images from the user study for personalized multi-human scene generation.
Figure S25: Example images from the user study for pose-controlled multi-human scene generation.
@@ B.3 Details on Baselines
To evaluate PersonaCraft, we compared it with several baselines for single-shot, multi-identity, and pose-controllable image synthesis, all implemented using SDXL [64]. Key baselines include OMG [43] for multi-concept personalization, InstantID/IPAdapter [81, 90] for single-shot personalization, 2D pose ControlNet [91], and optimization-based methods like DreamBooth [66] and Texture Inversion [22].
OMG + InstantID/IPAdapter/IPA-Face. OMG [43] introduces a two-stage sampling solution for multi-concept personalization. The first stage handles layout generation and occlusion management, while the second stage integrates concepts using visual comprehension and noise blending. OMG can also be combined with single-concept models like InstantID without additional tuning. For OMG+InstantID, we follow the official inference code from the InstantID repository [44]. For OMG+InstantID and OMG+IPAdapter/IPA-Face, we replace InstantID with IPAdapter and IPA-Face, respectively, to adapt the framework for different face identity modules.
@@ B.4 Details on User study
Personalized Multi-Human Scene Generation We conducted a user study to assess the naturalness, face identity preservation, body shape preservation, and text-image correspondence of images generated by three baseline methods (one from each group) and our method. Participants ranked the top three methods based on the following criteria: 1) Text Correspondence: Rank the images based on how closely they align with the textual description. 2) Face Identity Preservation: Rank the images in order of how well they reflect the face identity of the personalized character. 3) Body Shape Personalization: Rank the images based on how accurately they reflect the personalized character’s body shape. 4) Naturalness: Rank the images in order of the most natural-looking, considering factors such as physically impossible appearances, illogical features, inconsistencies, or lack of real-world physics and connections. The study collected a total of 18,540 responses from 103 participants across 15 cases, including both custom and COCO-Wholebody scenarios. We present illustrative example images from the user study in Fig. S24.
Pose-Controlled Multi-Human Scene Generation We conducted a user study to assess the naturalness, face identity preservation, body shape preservation, and text-image correspondence of images generated by three baseline methods (one from each group) and our method. Participants ranked the top three methods based on the following criteria:
1) Text Correspondence: Rank the images based on how closely they align with the textual description. 2) Pose Consistency: Rank the images based on how well they reflect the given pose input. 3) Naturalness: Rank the images in order of the most natural-looking, considering factors such as physically impossible appearances, illogical features, inconsistencies, or lack of real-world physics and connections. The study collected a total of 15,390 responses from 114 participants across 15 cases, including both custom and COCO-Wholebody scenarios. We present illustrative example images from the user study in Fig. S25.